# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v2

@@ III-D Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the $top-K$ predictions - commonly known as Kendall’s Tau [28]. For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU [67]), Precision, Recall, and F1 Score [15] given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU [45] score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. As mentioned earlier, if a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves.
IV Related Work
Conventional compiler testing is a domain that has been explored [10, 73, 16], involving differential testing [48, 65], as well as including compiler optimizations [11, 27, 19]. However, in terms of testing AI tools and compilers, most of the existing research has focused on (1) the direction of compiler fuzzing and (2) examining context-specific AI tools and APIs, such as DL framework converters.
In terms of compiler fuzzing, HirGen [43] is a utility for Deep Learning compiler fuzzing, focusing on the high-level model intermediate representation (IR) of TVM. TVMFuzz [1] allows generic-purpose TIR fuzzing on Apache TVM, detecting primarily crash and bit manipulation bugs. Tzer [38] is a coverage-guided fuzzer that applies modifications on models compiled in Apache TVM utilizing low-level IR (Tensor IR). GenCoG [69] focuses on generating valid and expressive TVM computation graphs by defining a domain specific language for constraint definition, as well as concolic constraint solving. NNSmith [37] focuses on valid DNN graph generation, usable to examine the compiler for bugs similarly with traditional grammar-based fuzzing techniques (e.g., [73, 16]), and MLIRSmith [68] integrates NNSmith’s fuzzing capabilities in MLIR [32] randomly. Suo et al. [60] also follow an operation dependency approach to guide the fuzzing process in MLIR. DeepDiffer [35] applies low-level IR fuzzing in a coverage-guided manner and comparing results across different DL compilers. SYNTHFUZZ [34] combines grammar-based fuzzing with custom mutation synthesis in order to effectively fuzz MLIR dialects. Li et al. [33] propose a methodology to uncover hardware-related faults by generating fault-sensitive, adversarial tests. Finally, Su et al. [59] fuzz DL compilers in the context of dynamic features, detecting bugs in PyTorch and its underlying tensor compiler.
Regarding testing techniques for context-specific AI tools and APIs, CRADLE [47] performs an execution analysis on the model graph in order to detect faults, while LEMON [70] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model. FAME [57] is based on LEMON but enables API mutation and optimizations for layer and weight mutation, guided by the detection of inconcistency faults. Similarly, Audee [20] aims to detect logical, not-a-number bugs, and crashes by applying an exploratory approach in combination with mutation testing. Finally, Louloudakis et al. propose FetaFix [42] - a utility that allows automatic fault detection in model converters, involving ONNX as an intermediate model representation.
Overall, limited work has been done in the direction of testing AI compiler optimizations. Choi et al. [14] have conducted an empirical study towards optimizer comparisons, however they do not consider graph-based optimizations in the post-training phase such as the ONNX Optimizer) as they focus on stochastic optimizers focusing primarily on the model training process, such as Adam [29]. DiffChaser [72] explores the application of differential testing across optimized and unoptimized DNN model versions. Louloudakis et al. propose DeltaNN [41], a utility to explore model robustness across different configurations - including but not limited to compiler optimizations. Xiao et al. [71] propose a metamorphic testing methodology for AI compiler bug detection, including issues related to optimizations.
To the best of our knowledge, OODTE is the first that thoroughly tests the ONNX Optimizer (and AI optimizers in general), using differential testing in order to detect faults. The closest existing work to our methodology, is DiffChaser [72], DeltaNN [41] and FetaFix [42]. However, these tools considerably differentiate from our contribution. In particular, DiffChaser attempts to detect problems related to AI optimization techniques (such as quantization) related to the models under test, not exploring potential faults introduced from bugs in the optimizer. DeltaNN applies differential testing, but focuses on uncovering behavioral inconsistencies in the models under test when deployed using different configurations, rather than focusing on compiler issues. FetaFix, on the other hand, aims to uncover issues related to converters using differential testing and involving ONNX. Its approach however does not consider model optimizations or AI optimizer tools.
V Experiment Setup
V-1 The Experiment Process
@@
We evaluated all ONNX Optimizer passes (as detailed in Section II-C). For each original model under test, we first applied the default setting of the ONNX Optimizer, which applies all fusion and elimination passes at once, generating the optimized model. We then compared the outputs of the original and optimized models according to the model type, as described in Section III-D. For optimized models presenting output differences, we performed optimization using only a single optimization pass, resulting in $47$ such single-pass optimized models. To identify which pass causes the problematic behavior, we compared the outputs of each single-pass model against those of the original model. It is worth noting that OODTE is compatible with new or custom passes as well. Furthermore, our methodology generalizes to other optimization frameworks (e.g., Apache TVM), provided they support both bundled and individual pass applications.
V-2 DNN Models
@@
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub [7]. In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet [24], EfficientNet-Lite4 [63] DenseNet121 [25], and InceptionV2 [61]), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet [36], Mask/Faster R-CNN R-50-FPN [21, 54], and SSD [39]), and $7$ transformer models [66] utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD [18]), sentiment analysis (RoBERTa [40]), text summarization (T5 [50]) and generic-purpose text generation (GPT-2 [49]), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.11.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
V-3 Computational Resources
@@
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API [2], or an AI compiler such as Apache TVM [12]).
V-4 Datasets
@@
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 [55] validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models. For text comprehension and generation, we used SQuAD-v2 [51], which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset [44, 64], containing $25000$ entries.
@@
    For image recognition models performing classification tasks, we utilized Kendall Rank Correlation Coefficient [28] (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be $<1$ for each input, we counted this input as different.
  2. 2.
@@
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different Intersection-over-Union (IoU) thresholds ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. 3.
@@
    For text generation models performing question answering and summarization, we utilized BLEU score [45] across output texts for each input, monitoring instances where $BLEU\_Score<1$ for each input.
  4. 4.
@@
    For sentiment analysis models, we compared each output value consisting of ${0,1}$ across the original and the optimized versions, registering any differences observed.
VI Results
@@
Following the execution of the full set of experiments11footnotemark: 1, OODTE successfully identified $15$ bugs, reported across $11$ separate GitHub issues [9]. A complete list of these issues, along with the associated models, is shown in Table I. Of the 15 bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
@@
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_to_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.

### v3

@@ 3.4. Metrics Comparison Module
Once the run completes per-chunk for both the original and the optimized model, the results are propagated to the Metrics Comparison Module. This module performs comparison against the results, by selecting the appropriate comparator based on the model type. For classification types, OODTE supports first-label comparison, but also calculates the Kendall Rank Correlation Coefficient across the topKtop-K predictions - commonly known as Kendall’s Tau (kendall, ). For Object Detection models, it utilizes the same methods of classification for labels and ranks, but also applies Intersection over Union (IoU (IoUBoundingBox, )), Precision, Recall, and F1 Score (F1Score, ) given an IoU threshold. For text generation models (e.g., GPT-2), it utilizes BLEU (BLEU, ) score in order to compare the text generated out both the original and the optimized model, as well as direct tensor comparison for cases of question answering (Q&A). In addition OODTE is build in a manner that new comparators can be added, in cases that a new category of models is tested.
3.5. Fault Localization
Once a method is selected, the Metrics Comparison Module calculates the metrics based on model type and generates a report with all results, in JSON format. If a difference is detected across the original and the optimized model, then OODTE retrofits the reported data to the Optimizer Module. Then, the Optimizer Module generates model optimizations per-pass, reporting how each pass behaves. OODTE applies a fault localization methodology for three cases: when the optimizer crashes, when it generates an optimized model in an invalid form, and when the optimization succeeds but produces results inconsistent with the original model. In addition, OODTE registers any warnings that can indicate the presence of a potential performance fault (e.g., redundant initializers in the model). Upon detecting an issue or a problematic behavior, OODTE attempts to identify the responsible pass by generating per-pass optimizations, executing inference on the resulting models, and comparing their outputs with those of the original model. This process is applied iteratively for each of the passes supported by the optimizer, until the pass responsible for the detected issue is identified.
4. Related Work
In this work, we compare our work with contributions related to AI compilers, focusing on the following key aspects: (1) the ability to localize faults associated with graph optimization passes, (2) the detection of compiler bugs affecting The types of bugs considered are the ones resulting in (a) model accuracy (inference results), (b) compiler crashes, (c) successful compilations, which, however, produce malformed or incorrect models that fail to run (e.g., models with incorrect input nodes or parameters), and (d) issues impacting execution times, (3) the capability to localize faults in the ONNX domain, particularly within ONNX Optimizer, and (4) the detection of faults in graph optimization passes. Table 1 summarizes the features supported by each contribution. For each tool, Table 1 shows: (1) the underlying technique, (2) support for bug detection in ONNX, (3) the ability to localize faults (Flt Loc.), (4) the types of bugs detected and the corresponding graph optimization passes involved (Gr. Opt. Passes), (5) detection of label accuracy faults (Acc.), (6) detection of crash faults (Cr.), (7) detection of malformed models produced despite successful compilation (Malf.), (8) detection of bugs affecting execution times (Ex. T.), and (9) the compilers supported/evaluated.
Table 1. Comparison of features between OODTE and related work. Each technique is compared across the following parameters: testing of ONNX (ONNX?), ability for fault localization (Flt. Loc.), bug types, graph optimization passes (Gr. Opt. Passes), detection of accuracy changes (Acc.), crashes (Cr.), generation of malformed models due to miscompilation (Malf.), and issues affecting model execution times(T.) bugs, as well as the compilers explored on each contribution.
:*: Limited support of feature, $\dagger:$ Does not consider ONNX Optimizer or compile-time optimizations associated with ONNX, $\ddagger$}: Uses ONNX as an intermediate representation, but does not perform fault detection associated to it.
Tool Technique ON NX? Flt. Loc. Bug Types Gr. Opt. Passes Acc. Cr. Malf. T. Compilers
OODTE
Per-pass fault
localization,
differential
testing
Logic, API versioning,
graph input/parameters/
transformations
Fuse/
eliminate/
rewrite
ONNX
Optimizer (onnxoptimizer, )
NNSmith (NNSmith, )
Grammar-
based fuzzing
†,‡ $\times$
Conversions, semantics,
gen.transformations,
graph rewrites
Conv2D
inputs
rewrite
$\times$ $\times$
TVM (tvm, ),
PyTorch JIT (pytorch, ),
TensorRT (tensorrt, ),
PT2 (pytorch2, ), ONNX
Runtime (onnxruntime, ),
TF XLA (tfxla, ),
TFLite (tensorflow, )
MLIRSmith (MLIRSmith, )
Grammar-
based fuzzing
$\times$ $\times$
Conversions (dialects),
generic transformations,
logic (dialects)
$\times$ $\times$ $\times$ $\times$
MLIR (MLIR, ),
LLVM (LLVM, )
MLIRod (SuoMLIR, )
Operator
dependency
analysis/
fuzzing
$\times$ $\times$
Conversions (dialects),
gen. transformations
$\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
SYNTH-
FUZZ (synthfuzz, )
Grammar-
based fuzzing
$\times$ $\times$ Generic, input-based $\times$ $\times$ $\times$ $\times$ MLIR (MLIR, )
HirGen (HirGen, )
Graph IR
fuzzing
$\times$ $\times$
Numeric computations,
exceptions, types, memory
$\times$ $\times$ $\times$ $\times$ TVM (tvm, )
TVMFuzz (TVMFuzz, )
Low-level
IR fuzzing
$\times$ $\times$
Numeric computations,
graph rewrites
Constant
folding
$\times$ $\times$ TVM (tvm, )
Tzer (TZER, )
Coverage-
guided fuzzing
(Low-level IR)
$\times$ $\times$
API inconsistency/misuse,
memory, types
$\times$ $\times$ TVM (tvm, )
GenCoG (GenCoG, )
Constraint
solving,
High-level
IR Fuzzing
$\times$ $\times$
Memory, tensor shape
/data, operator attributes,
logic
Operator
fusion
$\times$ $\times$ TVM (tvm, )
DeepDiffer (DeepDiffer, )
Low-level IR/
coverage-
guided fuzzing,
differential
testing
$\times$ $\times$
Type, memory, API misuse,
numeric computations,
exception handling
$\times$ $\times$ $\times$ TVM (tvm, )
WhiteFox (WhiteFox, )
LM analysis,
fuzzing
$\times$ $\times$
Memory, API misuse,
concurrency, logic,
security vulnerabilities
$\times$ $\times$
TF XLA (tfxla, ),
PyTorch
Inductor (pytorch, ),
LLVM (LLVM, )
TorchProbe (Su_2023, )
Feature-
based fuzzing
$\times$ $\times$
Memory, API misuse,
logic
$\times$ $\times$ $\times$
PyTorch (pytorch, ),
Triton (TilletTriton, )
MT-
DLComp (Xiao_2022, )
Metamorphic
testing
$\times$ $\times$
Memory, dead code,
layout optimization,
numeric accuracy
Conv2D
layout
change
$\times$ $\times$
TVM (tvm, ),
Glow (Glow, ),
NNFusion (NNFusion, ),
TF XLA (tfxla, )
FetaFix (FetaFix, )
Fault localiza-
tion,
differential
testing
$\times$
Preprocessing, input
dimensions, parameters,
hyperparameters, tensor
shape, graph structure
$\times$ $\times$ $\times$ $\times$
Converters across
PyTorch (pytorch, ),
Keras (chollet2015keras, ),
TF (tensorflow, ), TFLite (tensorflow, )
In terms of the Technique column shown in Table 1, most existing works primarily rely on compiler fuzzing, often enhanced with techniques such as differential testing (e.g., DeepDiffer (DeepDiffer, )) or LLM analysis (e.g., WhiteFox (WhiteFox, )). Consequently, these systems focus on triggering compiler bugs rather than pinpointing their origins and root causes, typically relying on generic compiler error messages to perform manual fault localization. Among these works, only FetaFix (FetaFix, ) provides a mechanism for fault localization, but it targets accuracy-related faults in model conversion between DL frameworks and does not support the model optimization process. In contrast, OODTE systematically localizes faults by applying an iterative fault localization methodology for each pass in the ONNX Optimizer.
In terms of bug types detected in Table 1, existing tools primarily focus on detecting crash bugs. For label accuracy-related bugs, they typically rely on random input generation (e.g., NNSmith (NNSmith, )) to expose inconsistencies in model inference. In contrast, for label inconsistency bugs that OODTE detects, we assess the prevalence and impact of these bugs on popular datasets(e.g., ILSVRC (ILSVRC17, ), IMDB (imdbsentiment, )) and model architectures. Additionally, only a small number of tools consider miscompilations (for instance, models that contain problematic input nodes or misconfigured parameters), with just two prior works capable of detecting faults of this kind (TVMFuzz, ; WhiteFox, ). The detection of bugs affecting execution times of the compiled models has received limited attention in the literature, with only two TVM-focused systems (TZER, ; DeepDiffer, ) using empirical measurements to evaluate performance-related faults. OODTE currently identifies issues that could potentially lead to degraded model execution times by flagging redundant operator warnings (e.g., unused initializer entries), while the integration of precise execution time measurements is left as subject for future work.
Regarding support for ONNX in Table 1, OODTE is the only tool capable of localizing faults in the ONNX Optimizer. Other methodologies either treat ONNX merely as an intermediate representation or focus on bug detection in ONNX Runtime (WhiteFox, ), a cross-platform inference engine, unlike model optimization targeted by the ONNX Optimizer.
Existing work on bug detection within graph optimization passes (shown as Gr. Opt. Passes column in Table 1) is limited, with prior approaches focusing on common operations such as constant folding, operator fusion, and convolution optimizations. In contrast, OODTE addresses the full spectrum of features in the ONNX Optimizer, covering a set of 47 passes that perform a wide variety of graph optimizations such as Fuse, Eliminate, and Rewrite operations in addition to the more common operations.
5. Experiment Setup
5.1. The Experiment Process
@@
We evaluated all ONNX Optimizer passes (Section 2.3). For each model, we first applied the default setting, which runs all fusion and elimination passes, and compared the outputs of the original and optimized models as described in Section 3.4. When discrepancies arose, we re-optimized using individual passes, producing 47 single-pass models to identify the pass causing the issue. OODTE also supports new or custom passes, and our methodology generalizes to other optimization frameworks (e.g., Apache TVM) that allow both bundled and individual pass execution.
5.2. DNN Models
@@
For our experiment set, we sourced $130$ model instances from the ONNX Model Hub (onnxmodelhub, ). In particular, we selected $93$ models utilized for classification tasks (e.g., MobileNet (mobilenet, ), EfficientNet-Lite4 (efficientnet, ) DenseNet121 (densenet, ), and InceptionV2 (inception, )), $30$ models used for object detection and semantic segmentation tasks (e.g., RetinaNet (retinanet, ), Mask/Faster R-CNN R-50-FPN (mask-rcnn, ; faster-rcnn, ), and SSD (ssd, )), and $7$ transformer models (attention, ) utilized for text comprehension and generation tasks, such as question answering (Q&A), (BERT-SQuAD (BERT, )), sentiment analysis (RoBERTa (liu2020roberta, )), text summarization (T5 (T5Paper, )) and generic-purpose text generation (GPT-2 (gpt2, )), with $4$ models focusing on Q&A, while one model is utilized for each of the other text comprehension and generation tasks. We utilized models of a variety of opsets based on hub availability, ranging from $7$ (the lowest compatible with the latest ONNX Runtime version by the time of this publication - v1.21.11.21.1) up to $12$ (the highest provided from ONNX Model Hub via their API). Our methodology, however, is applicable to any ONNX model type and is not constrained by opset versions or tied to a specific model repository.
5.3. Computational Resources
@@
We ran our models on mid-range systems with Intel i5 and i7 CPUs, reflecting the typical hardware used by developers during model development and fine-tuning. Our method is generalizable to larger models, limited only by ONNX model size constraints, and can be extended to high-end systems with a trivial extension for hardware acceleration (using the ONNX Runtime API (onnxruntime, ), or an AI compiler such as Apache TVM (tvm, )).
5.4. Datasets
@@
We selected datasets appropriate to each model task. For classification and object detection, we used the ILSVRC2017 (ILSVRC17, ) validation dataset, containing $5500$ images, commonly used for benchmarking image recognition models (e.g., ResNet (resnet, ) and BERT (BERT, )). For text comprehension and generation, we used SQuAD-v2 (rajpurkar2018knowdontknowunanswerable, ), which includes $11873$ questions and contexts, to evaluate Q&A tasks and general-purpose transformer behavior. Additionally, for summarization and sentiment analysis, we used the IMDB dataset (moviedataset, ; imdbsentiment, ), containing $25000$ entries.
@@
    For image recognition models performing classification tasks, we utilized the Kendall Rank Correlation Coefficient (kendall, ) (KT), measuring differences for top-K prediction labels of each dataset input image, for $K=1,5$, and $10$. If the KT value was found to be <1<1 for each input, we counted this input as different.
  2. (2)
@@
    For image recognition models performing object detection and semantic segmentation, we calculated precision, recall, and F1 score for each dataset input, and extracted their average, combined with different thresholds of Intersection-over-Union (IoU) ($IoU=0.5,0.75$, and $0.9$), as well as IoU, mean Average Precision (mAP), along with average precision and recall. We considered the original model output as the base, and the optimizer model the one under test for our metric calculations, and the higher the values observed, the better the optimized model aligns with the original.
  3. (3)
@@
    For text generation models performing question answering and summarization, we utilized BLEU score (BLEU, ) across output texts for each input, monitoring instances where BLEU_Score<1BLEU\_Score<1 for each input.
  4. (4)
@@
    For sentiment analysis models, we compared each output value consisting of 0,1{0,1} across the original and the optimized versions, registering any differences observed.
6. Results
@@
Following the execution of the full set of experiments. A complete list of these issues, along with the associated models, is shown in Table 2. The developers have acknowledged the issues we reported, and we have provided them with full reproducibility details for each reported bug. Of the $15$ bugs identified, $14$ were entirely new (not previously reported on the ONNX Optimizer issue tracker), while one was a previously reported issue that lacked sufficient detail, including model information and reproducibility steps. We provided comprehensive information and full reproducibility details for this issue.
@@
Within the ONNX optimizer, the bugs identified were primarily associated with $9$ of the $47$ passes. However, not all passes are equally important. By default, the optimizer prioritizes passes related to fuse and eliminate operations. As a result, $8$ of the $11$ reported GitHub issues are considered more critical since they belong to the optimizer’s core (default) strategy. Among these $8$ issues, $6$ are linked to specific passes (e.g., fuse_bn_into_conv, which fuses batch normalization nodes with convolutions), while the remaining 2 issues were more general and related to the optimizer’s overall usage. We analyze and present the impact of these faults both on the optimization process and on the correctness of the optimized models in the following Sections.
6.1. Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes (1717%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN (faster-rcnn, ), RetinaNet (retinanet, ), VGG-16, VGG-19 (vgg, ), as well as YOLOv3 (Redmon2018YOLOv3AI, ), presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances (9.29.2%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
As shown in Table 2, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases). Overall, object detection models were more prone to optimizer crashes compared to classification models. Due to their increased structural complexity and larger size, they are more likely to exhibit errors related to node properties and initializers, which account for the vast majority of crash faults. In contrast, only one issue was observed in the most complex model category — text generation models, that was related to model input in GPT-2. However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.